A firm borrows P2,000 for 6 years at 8%. At the end of 6 years, it renews the loan for the amount due plus P2,000 more for 2 years at 8%. What is the lump sum due?

1
Expert's answer
2022-01-31T14:22:13-0500

"A1=6000(0.0067+\frac{0.0067}{(1+0.0067)^{72}-1})=104.79"

"104.79\times72=7545,075"


7545,075+2000=9545,075


"A2=9545.075(0.0067+\frac{0.0067}{(1+0.0067)^{24}-1})=429.7234"


"429.7234\times24=10313.36"